OpenStudy (anonymous):

SOMEONE PLEASE HELP!! ILL MEDAL AND FAN! Two equations are shown: Equation 1: 3 4 (x−12)=12 Equation 2: 3 4 y−12=12 Solve each equation. Then, enter a number in each box to make this statement true. The value of x is: The value of y is:

OpenStudy (anonymous):

@Susanna @foolofatook @annie12m Can any of you please help?

OpenStudy (mathmale):

Please clarify the problem. I don't understand what the 3 4 represents in your 3 4 (x−12)=12.

OpenStudy (anonymous):

it means 3/4 or 3 over 4 @mathmale

OpenStudy (kurayami95):

are you meaning that the equation is 3/4(x-12)=12 ? or the 3/4 is apart ?

OpenStudy (anonymous):

The equation is 3/4(x-12)=12 @Kurayami95

OpenStudy (kurayami95):

oh then you should distribute first

OpenStudy (mathmale):

sorry to seem picky, but it's essential that you learn how to present math problems accurately here. Otherwise, time and effort (including yours) will be lost. In this case, I'd suggest you write "three fourths" as (3/4), or \[\frac{ 3 }{ 4 }, \]

OpenStudy (mathmale):

or 

OpenStudy (mathmale):

By 3 4 (x−12)=12 did you mean\[\frac{ 3 }{ 4 }(x-12)=12?\]

OpenStudy (kurayami95):

she did mean that i believe so

OpenStudy (mathmale):

If I've gotten that wrong, please fix it. Else, focus now on the 2nd equation and put it into proper form.

OpenStudy (anonymous):

My internet isnt working good, it keeps shutting on and off sorry for not replying so quickly, but yes that is what i meant to say @mathmale

OpenStudy (mathmale):

So, your first equation is \[\frac{ 3 }{ 4 }(x-12)=12\]and your second is ... ?

OpenStudy (mathmale):

Surprisingly, these two equations are independent...we don't need one to solve the other. All you need to do with the first is to "solve for x."

OpenStudy (mathmale):

I'll wait for you to post your second equation here.

OpenStudy (anonymous):